# Online estimation of the inverse of the Hessian for stochastic optimization with application to universal stochastic Newton algorithms

### v1

@@ 6.1 Study on the largest eigenvalue of $A_{n}$
$=A_{n}-H^{-1}-\gamma_{n+1}\left(\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}\right)+\gamma_{n+1}\xi_{n+1}+$
$\quad\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}$
where $\xi_{n+1}:=-W_{n+1}+\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})$. Let $\alpha^{*}_{k}$ be the function defined for all $h\in\mathcal{M}_{p}(\mathbb{R})$ by
@@
$A_{n+1}-A$ $=\alpha_{n}^{*}\left(A_{n}-H^{-1}\right)-\gamma_{n+1}^{2}H(A_{n}-H^{-1})H$
$+\gamma_{n+1}r_{1,n}+\gamma_{n+1}r_{2,n}+\gamma_{n+1}s_{n}+\gamma_{n+1}s^{\prime}_{n}+\gamma_{n+1}\xi_{n+1}$ (16)
@@
$+\left(\nabla^{2}G(\hat{\theta}_{n})-H\right)\left(A_{n}-H^{-1}\right)+\left(A_{n}-H^{-1}\right)\left(\nabla^{2}G(\hat{\theta}_{n})-H\right),$
@@
$\quad+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}r_{2,k}+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s_{k}+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s^{\prime}_{k},$ (17)
@@
Next, we will determine the rate of convergence for each term on the right-hand side of equation (6.2).
Rate of convergence of $M_{n}:=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}\xi_{k+1}$: Recall that for all $\delta>0$, $\lambda_{\max}(A_{n})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.$ Thus, there exists a constant $c^{\prime}>0$ such that $\mathbb{E}\left[\left\|\xi_{n+1}\right\|^{2}_{F}|\mathcal{F}_{n}\right]\leq c^{\prime}\left(1+\left\|A_{n}\right\|_{F}^{2}\right)=o\left(n^{2-2\gamma}\ln n^{2+2\delta}\right)$ a.s., and according to Lemma 6.3, it follows that
@@
$\leq\left(1-\lambda_{\min}(H)\gamma_{n+1}\right)^{2}\left\|R_{1,n}\right\|_{op}+\gamma_{n+1}\left\|r_{1,n+1}\right\|_{op}a.s.$
By Assumption (A4), $\left\|r_{1,n+1}\right\|_{op}=o\left(n^{-a/2}\ln n^{(1+\delta)/2}\right)a.s.$ According to Lemma 6.4,
@@
Rate of convergence of $R_{2,n}:=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}r_{2,k}$:
@@
$\leq$ $12\mathbb{E}\left[\left(\left\|Q_{n+1}\right\|^{2}\left\|Z_{n+1}\right\|^{2}\left\|A_{n}\right\|^{2}+d\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right].$
Applying Markov’s inequality, and since $\left\|Z_{n}\right\|$ is supposed to be bounded by $M$,
$\mathbb{E}\left[\left\|(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}\right\|^{2}|\mathcal{F}_{n}\right]$ $\leq 12d\mathbb{E}\left[\left\|\nabla^{2}g(X_{n+1},\hat{\theta}_{n})\right\|_{op}^{q}\left\|Z_{n}Z_{n}^{T}\right\|^{q}|\mathcal{F}_{n}\right]\beta_{n}^{-q}$
$+12\left\|A_{n}\right\|^{2}\mathbb{E}\left[\left\|Q_{n}\right\|^{2}M^{2}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}\right]$
$\leq 12dM^{2q}\mathbb{E}\left[\left\|\nabla^{2}g(X_{n+1},\hat{\theta}_{n})\right\|_{op}^{q}|\mathcal{F}_{n}\right]\beta_{n}^{-q}$
$+12C_{q}\left\|A_{n}\right\|^{2}\beta_{n+1}^{2-q}$
Thus, by Lemma 6.4,
@@
$+2q\mathbb{E}\left[\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right]=\mathcal{O}\left(\beta_{n}^{1-q}\right).$
Therefore, thanks to Lemma 6.4,
@@
A first result for $A_{n}$: Thanks to equalities (18) to (21), one can rewrite $A_{n}-H^{-1}$ as
$A_{n}-H^{-1}=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s_{k}+\widetilde{S}_{n}$
with $\left\|\widetilde{S}_{n}\right\|_{F}=o\left(n^{\max\left(1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\right)}\ln n^{1+\delta}\right).$ In addition, we have
@@
$\quad+\left\|\nabla^{2}G(\hat{\theta}_{n})-H\right\|\left\|A_{n}-H^{-1}\right\|_{op}$
@@
$\left\|A_{n}-H^{-1}\right\|_{F}=o\left(n^{\max\left(1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\right)}\ln n^{1+\delta}\right)a.s.$
Final rate of convergence of $A_{n}$: The aim here is, with the help of this first result on $A_{n}$, to give better rates of convergence for $M_{n}$ and $R_{2,n}$. First, note that if $\gamma>2/3$, then $01-\frac{3}{2}\gamma<$ 0 and we have directly $A_{n}\xrightarrow[n\rightarrow\infty]{a.s.}H^{-1}$. If $\gamma\leq 2/3$, we have $01-\frac{3}{2}\gamma>$ 0 , so that $\max\{1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\}=\max\{1-\frac{3}{2}\gamma,\beta(1-q)\}$ and $\left\|A_{n}\right\|=o\left(n^{\max\{1-\frac{3}{2}\gamma,\beta(1-q)\}}\ln n^{1+\delta}\right)a.s.$. Thus, applying Lemma 6.3 in the worse case (i.e when $\gamma>2/3$), one now has
@@
Following the same process as before, we now have that for any $\gamma\in(1/2,1)$, $A_{n}$ converges almost surely to $H^{-1}$, and in particular, one has $\left\|A_{n}\right\|$ = $\mathcal{O}\left(1\right)$. Applying another time Lemma 6.3, we have
@@
Rate of convergence of $A_{n,\tau}$. Decomposition (16) can be written as
$H(A_{n}-H^{-1})+(A_{n}-H^{-1})H=\frac{A_{n}-H^{-1}-(A_{n+1}-H^{-1})}{\gamma_{n+1}}+r_{1,n}+r_{2,n}+s_{n}+s^{\prime}_{n}+\xi_{n+1},$
@@
$H(A_{n,\tau}-H^{-1})+(A_{n,\tau}-H^{-1})H=$ $t_{n}\sum_{k=0}^{n}u_{k}\left((A_{k}-H^{-1})-(A_{k+1}-H^{-1})\right)$
$+t_{n}\sum_{k=0}^{n}u_{k}\gamma_{k+1}(r_{1,n}+r_{2,n}+s_{n}+s^{\prime}_{n}+\xi_{n+1})$ (23)
where $t_{n}=\frac{1}{\sum^{n}_{k=0}\log(k+1)^{\tau}}$ and $u_{k+1}=\frac{\ln(k+1)^{\tau}}{\gamma_{k+1}}$. Our objective is to determine the rate of convergence for each term on the right-hand side of the decomposition given in (6.2).
@@
$t_{n}\left\|A_{n+1}-H^{-1}\right\|u_{n+1}=o\left(\frac{\ln n^{1+\delta}}{n^{1-\gamma+\frac{1}{2}\min\left\{\gamma,a,2\beta(q-1)\right\}}}\right)\quad a.s$
@@
$\left\|t_{n}\sum_{k=0}^{n}u_{k}\left((A_{k}-H^{-1})-(A_{k+1}-H^{-1})\right)\right\|_{F}=o\left(\frac{\ln n^{1+\delta}}{n^{1-\gamma+\frac{1}{2}\min\left\{\gamma,a,2\beta(q-1)\right\}}}\right)\quad a.s$
Rate of convergence of $t_{n}\sum_{k=0}^{n}u_{k+1}\gamma_{k+1}\xi_{k+1}$. Remark that there exists $c^{\prime}>0$ such that for all $\delta>0$
$\mathbb{E}\left[\left\|\xi_{n+1}\right\|^{2}_{F}\|\mathcal{F}_{n}\right]\leq c^{\prime}\left(1+\left\|A_{n}\right\|_{F}^{2}\right)\leq c^{\prime}\left(1+2\left\|A_{n}-H^{-1}\right\|_{F}^{2}+2\left\|H^{-1}\right\|_{F}^{2}\right).$
@@
Then, recalling that $\left\|r_{1,k+1}\right\|_{op}=o\left(n^{-a/2}\ln k^{(1+\delta)/2}\right)$
@@
Rate of convergence of $t_{n}\sum_{k=0}^{n}u_{k+1}\gamma_{k+1}s_{k}^{\prime}$. Recalling that $s_{k}^{\prime}=O\left(\beta_{k}^{1-q}\right)$, it comes
@@
$\quad+\left\|\nabla^{2}G(\hat{\theta}_{n})-H\right\|\left\|A_{n}-H^{-1}\right\|_{op},$
@@
$\left\|A_{n}-H^{-1}\right\|_{op}\mathbb{E}\left[\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n}}\|\mathcal{F}_{n}\right]$
@@
leading to
@@
$\left\|t_{n}\sum_{k=0}^{n}u_{k}\gamma_{k+1}s_{k}\right\|_{F}^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\min\{2,2\min\left\{a,\gamma,2\beta(q-1)\right\}\}}}\right)a.s.$
@@
Note that $H$ is diagonalisable, and its eigenvalues are denoted by $\lambda_{1},...,\lambda_{d}$. Thus, for any matrix $B\in\mathcal{M}_{d}$, we have
@@
In order to prove this theorem, we first show that the eigenvalues of $A_{n}$ and $A_{n,\tau}$ are well controlled, which implies the consistency of $\theta_{n}$. Then, we give an initial rate of convergence of $\theta_{n}$, which allows us to obtain the rate of convergence of $A_{n}$ and $A_{n,\tau}$. The last step is to deduce the rate of convergence of $\theta_{n}$ and $\theta_{n,\tau}$. We first study the smallest eigenvalue of $A_{n}$.
@@
Under Assumptions(A3) and (A5), the smallest eigenvalue of $A_{n}$ denoted by $\lambda_{\min}(A_{n})$ satisfies

### v3

@@ 6.4.1 Rate of convergence of $A_{n}$
$=A_{n}-H^{-1}-\gamma_{n+1}\left(\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}\right)+\gamma_{n+1}\xi_{n+1}{+\Pi_{n+1}^{\perp}}$
where $\xi_{n+1}:=-W_{n+1}+\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})$. Let $\alpha^{*}_{k}$ be the function defined for all $h\in\mathcal{M}_{d}(\mathbb{R})$ by
@@
$A_{n+1}-H^{-1}=\alpha_{n}^{*}\left(A_{n}-H^{-1}\right){+\Pi_{n+1}^{\perp}}$ $+\gamma_{n+1}r_{1,n}+\gamma_{n+1}r_{2,n}+\gamma_{n+1}s_{n}+\gamma_{n+1}s^{\prime}_{n}+\gamma_{n+1}\xi_{n+1}$ (16)
@@
$+\left(\nabla^{2}G(\hat{\theta}_{n})-H\right)\left(A_{n}-H^{-1}\right)+\left(A_{n}-H^{-1}\right)\left(\nabla^{2}G(\hat{\theta}_{n})-H\right)-\gamma_{n+1}^{2}H(A_{n}-H^{-1})H,$
@@
$\quad+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}r_{2,k}+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s_{k}+\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s^{\prime}_{k}$
$\quad{+\sum_{k=0}^{n-1}\Psi_{n,k+1}^{*}\gamma_{k+1}\Pi_{k+1}^{\perp}}$ (17)
@@
Next, we will determine the rate of convergence for each term on the right-hand side of equation (6.4.1).
Rate of convergence of $M_{n}:=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}\xi_{k+1}$: Recall that for all $\delta>0$, $\lambda_{\max}(A_{n})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.$ Thus, there exists a constant $c^{\prime}>0$ such that $\mathbb{E}\left[\left\|\xi_{n+1}\right\|^{2}_{F}|\mathcal{F}_{n}\right]\leq c^{\prime}\left(1+\left\|A_{n}\right\|_{F}^{2}\right)=o\left(n^{2-2\gamma}\ln n^{2+2\delta}\right)$ a.s., and according to Lemma B.3, it follows that
@@
$\leq\left(1-\lambda_{\min}(H)\gamma_{n+1}\right)^{2}\left\|R_{1,n}\right\|_{op}+\gamma_{n+1}\left\|r_{1,n+1}\right\|_{op}$
By Assumption (A3),
@@
According to Lemma B.4,
Rate of convergence of $R_{2,n}:=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}r_{2,k}$: Observe that $\left(r_{2,n}\right)$ is a sequence of martingale differences adapted to the filtration $\mathcal{F}_{n}$. Moreover, since $\|A+B\|_{F}^{2}\leq 2\|A\|_{F}^{2}+2\|B\|_{F}^{2}$ (used twice),
@@
$\leq$ $8\mathbb{E}\left[\left(\left\|Q_{n+1}\right\|^{2}\left\|Z_{n+1}\right\|^{2}\left\|A_{n}\right\|_{F}^{2}+d\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right].$
Applying Markov’s inequality, and since $\left\|Z_{n+1}\right\|\leq M$,
$\mathbb{E}\left[\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right]\leq\beta_{n+1}^{-q}\mathbb{E}\left[\left\|\nabla_{h}^{2}g\left(X_{n+1},\hat{\theta}_{n}\right)\right\|_{op}^{q}\left\|Z_{n+1}\right\|^{2q}|\mathcal{F}_{n}\right]\leq C_{q}M^{2q}\beta_{n+1}^{-q}.$
In a same way,
$\mathbb{E}\left[\left\|Q_{n+1}\right\|^{2}\left\|Z_{n+1}\right\|^{2}\left\|A_{n}\right\|^{2}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right]$ $\leq\beta_{n+1}^{2-q}\left\|A_{n}\right\|_{F}^{2}\mathbb{E}\left[\left\|Q_{n+1}\right\|^{q}\left\|Z_{n+1}\right\|^{q}|\mathcal{F}_{n}\right]$
$\leq C_{q}\left\|A_{n}\right\|_{F}^{2}M^{2q}\beta_{n+1}^{2-q}.$
Then,
$\mathbb{E}\left[\left\|(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}\right\|^{2}|\mathcal{F}_{n}\right]$ $\leq 8C_{q}M^{2q}\beta_{n}^{-q}+8C_{q}\left\|A_{n}\right\|_{F}^{2}M^{2q}$
Thus, with the help of Lemma B.3,
@@
$+2q\mathbb{E}\left[\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}|\mathcal{F}_{n}\right]$
$=\mathcal{O}\left(\beta_{n+1}^{1-q}\right).$ (22)
Therefore, thanks to Lemma B.4,
@@
Rate of convergence of $R_{\Pi,n}=\sum_{k=0}^{n-1}\Psi_{n,k+1}^{*}\gamma_{k+1}\Pi_{k+1}^{\perp}$: Observe that $\Pi_{n+1}^{\perp}$ is different from 00 if
$\left\|A_{n}-\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right\|_{F}\leq\beta_{n+1}^{\prime}.$
Nevertheless, one has
$\left\|A_{n}-\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right\|_{F}$ $\leq\left\|A_{n}\right\|_{F}\left(1+2\gamma_{n+1}\beta_{n+1}\right)+2d\gamma_{n+1}$
$=o\left(\ln(n+1)^{1+\delta}n^{1-\gamma}\right)\quad a.s.$
and since $\beta^{\prime}>1-\gamma$, it comes that $\mathbf{1}_{\Pi_{n+1}^{\perp}\neq 0}$ converges almost surely to 00. Then
$\|R_{\Pi,n}\|_{op}\leq\left\|\Psi_{n,0}^{*}\right\|_{op}\sum_{k=0}^{n-1}\left\|\Psi_{k,0}^{-1}\gamma_{k+1}\Pi_{k+1}^{\perp}\right\|_{op}\mathbf{1}_{\Pi_{k+1}^{\perp}\neq 0}=O\left(\left\|\Psi_{n,0}^{*}\right\|_{op}\right)\quad a.s.$
and this term converges exponentially fast to 00.
A first result for $A_{n}$: Thanks to equalities (18), (20), (21) and (23), one can rewrite $A_{n}-H^{-1}$ as
$A_{n}-H^{-1}=\sum_{k=0}^{n-1}\Psi^{*}_{n,k+1}\gamma_{k+1}s_{k}+\widetilde{S}_{n}+\Psi_{n,0}^{*}\left(A_{0}-H^{-1}\right)$
with $\left\|\widetilde{S}_{n}\right\|_{F}=o\left(n^{\max\left(1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\right)}\ln n^{1+\delta}\right)$ a.s. and $\Psi_{n,0}^{*}\left(A_{0}-H^{-1}\right)$ converges exponentially fast to 00. In addition, since $\hat{\theta}_{n}$ converges almost surely to $\theta$ and since
$\mathbb{E}\left[\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n}}\|\mathcal{F}_{n}\right]$ converges almost surely to 00 (one can perform analogous calculations to those used in deriving equality (20)),
@@
$\quad+\left\|\nabla^{2}G(\hat{\theta}_{n})-H\right\|\left\|A_{n}-H^{-1}\right\|_{op}+\gamma_{n+1}^{2}\|H\|_{op}^{2}\left\|A_{n}-H^{-1}\right\|_{op}$
@@
$\left\|A_{n}-H^{-1}\right\|_{F}=o\left(n^{\max\left(1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\right)}\ln n^{1+\delta}\right)\quad a.s.$ (24)
Final rate of convergence of $A_{n}$: The aim here is, with the help of this first result on $A_{n}$ given by (24), to obtain better rates of convergence for $M_{n}$ and $R_{2,n}$. First, note that if $\gamma>2/3$, then $01-\frac{3}{2}\gamma<$ 0 and we have directly $A_{n}\xrightarrow[n\rightarrow\infty]{a.s.}H^{-1}$. If $\gamma\leq 2/3$, we have $01-\frac{3}{2}\gamma>$ 0 , so that $\max\{1-\frac{3}{2}\gamma,-a/2,\beta(1-q)\}=\max\{1-\frac{3}{2}\gamma,\beta(1-q)\}$ and $\left\|A_{n}\right\|=o\left(n^{\max\{1-\frac{3}{2}\gamma,\beta(1-q)\}}\ln n^{1+\delta}\right)a.s.$. Thus, when $\gamma\leq 2/3$, applying Lemma B.3 and following the same calculus as for obtaining (18) and (21), one now has
@@
Following the same process as before, we now have that for any $\gamma\in(1/2,1)$, $A_{n}$ converges almost surely to $H^{-1}$, and in particular, one has $\left\|A_{n}\right\|$ = $\mathcal{O}\left(1\right)$ a.s. Applying another time Lemma B.3 and following the same calculus as for obtaining (18) and (21), we have
@@
6.4.2 Rate of convergence of $A_{n,\tau}$.
First, note that decomposition (16) can be written as
$H(A_{n}-H^{-1})+(A_{n}-H^{-1})H=\frac{A_{n}-H^{-1}-(A_{n+1}-H^{-1})}{\gamma_{n+1}}+r_{1,n}+r_{2,n}+s_{n}+s^{\prime}_{n}+\xi_{n+1}{+\frac{\Pi_{k+1}^{\perp}}{\gamma_{k+1}}},$
@@
$H(A_{n,\tau}-H^{-1})+(A_{n,\tau}-H^{-1})H=$ $t_{n}\sum_{k=0}^{n}u_{k+1}\left((A_{k}-H^{-1})-(A_{k+1}-H^{-1})\right)$
$+t_{n}\sum_{k=0}^{n}u_{k+1}\gamma_{k+1}\left(r_{1,k}+r_{2,k}+s_{k}+s^{\prime}_{n}+\xi_{k+1}{+\frac{\Pi_{k+1}^{\perp}}{\gamma_{k+1}}}\right)$ (26)
where $t_{n}=\frac{1}{\sum^{n}_{k=0}\log(k+1)^{\tau}}$ and $u_{k+1}=\frac{\ln(k+1)^{\tau}}{\gamma_{k+1}}$. Our objective is to determine the rate of convergence for each term on the right-hand side of decomposition (6.4.2).
@@
$t_{n}\left\|A_{n+1}-H^{-1}\right\|u_{n+1}=o\left(\frac{\ln n^{(1+\delta)/2}}{n^{1-\gamma+\frac{1}{2}\min\left\{\gamma,a,2\beta(q-1)\right\}}}\right)\quad a.s$
@@
$\left\|t_{n}\sum_{k=0}^{n}u_{k}\left((A_{k}-H^{-1})-(A_{k+1}-H^{-1})\right)\right\|_{F}=o\left(\frac{\ln n^{1+\delta}}{n^{1-\gamma+\frac{1}{2}\min\left\{\gamma,a,2\beta(q-1)\right\}}}\right)\quad a.s$ (27)
Rate of convergence of $t_{n}\sum_{k=0}^{n}u_{k+1}\gamma_{k+1}\xi_{k+1}$. Remark that there exists $c^{\prime}>0$ such that
$\mathbb{E}\left[\left\|\xi_{n+1}\right\|^{2}_{F}\|\mathcal{F}_{n}\right]\leq c^{\prime}\left(1+\left\|A_{n}\right\|_{F}^{2}\right)\leq c^{\prime}\left(1+2\left\|A_{n}-H^{-1}\right\|_{F}^{2}+2\left\|H^{-1}\right\|_{F}^{2}\right)=\mathcal{O}(1)\quad a.s.$
@@
Then, recalling that $\left\|r_{1,k+1}\right\|_{op}=o\left(n^{-a/2}\ln k^{(1+\delta)/2}\right)$ a.s. (see equality (19)),
@@
Rate of convergence of $t_{n}\sum_{k=0}^{n}u_{k+1}\gamma_{k+1}s_{k}^{\prime}$. Recalling that $s_{k}^{\prime}=O\left(\beta_{k}^{1-q}\right)$ (see equality (22)), it comes
@@
$\left\|s_{n}\right\|_{op}$ $\leq\left\|A_{n}-H^{-1}\right\|_{op}\mathbb{E}\left[\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n}}|\mathcal{F}_{n}\right]$
$\quad+\left\|\nabla^{2}G(\hat{\theta}_{n})-H\right\|\left\|A_{n}-H^{-1}\right\|_{op}+\gamma_{n+1}\|H\|_{op}\left\|A_{n}-H^{-1}\right\|_{op},$
@@
$\left\|A_{n}-H^{-1}\right\|_{op}\mathbb{E}\left[\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n}}|\mathcal{F}_{n}\right]$
@@
and the term $\gamma_{n+1}\|H\|_{op}\left\|A_{n}-H^{-1}\right\|_{op}$ is negligible, leading to
@@
$\left\|t_{n}\sum_{k=0}^{n}u_{k}\gamma_{k+1}s_{k}\right\|_{F}^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{2\min\left\{a,\gamma,2\beta(q-1)\right\}}}\right)a.s.$
@@
Rate of convergence of $t_{n}\sum_{k=0}^{n}\frac{\Pi_{k+1}^{\perp}}{\gamma_{k+1}}$: Since $\mathbf{1}_{\Pi_{k+1}^{\perp}\neq 0}$ converges almost surely to 00,
$\left\|t_{n}\sum_{k=0}^{n}\frac{\Pi_{k+1}^{\perp}}{\gamma_{k+1}}\right\|=O(t_{n})\quad a.s.$
and this term is so negligible.
@@
Note that $H$ is diagonalisable in an orthonormal basis, and its eigenvalues are denoted by $\lambda_{1},...,\lambda_{d}$. Thus, for any matrix $B\in\mathcal{M}_{d}$, we have
@@
In order to prove this theorem, we show that the eigenvalues of $A_{n}$ and $A_{n,\tau}$ are well controlled, which implies the consistency of $\theta_{n}$. Then, we give an initial rate of convergence of $\theta_{n}$, which allows us to obtain the rate of convergence of $A_{n}$ and $A_{n,\tau}$. The last step is to deduce the rate of convergence of $\theta_{n}$ and $\theta_{n,\tau}$.
@@
Under Assumptions(A3) and (A5), the smallest eigenvalue of $A_{n}$ (resp. $A_{n,\tau}$) denoted by $\lambda_{\min}(A_{n})$ (resp. $\lambda_{\min}(A_{n},\tau)$) satisfies